# Neural Context Flows for Meta-Learning of Dynamical Systems

### v4

@@ 3.2 Adaptation (or Meta-Testing)
Our adaptation rule, as outlined in Algorithm 2 is extremely fast, converging in seconds for trainings that took hours. In scenarios where we want to adapt to more than one environment, we outline a bulk version in Section A.5 (see Algorithm 4). Although this algorithm is parallelizable and framed in the same way as the meta-training algorithm, it does not allow for “flow” of contextual information, since this causes significant accuracy degradation. Most importantly, disabling the Taylor expansion at this stage limits the size of the context pool to $p=1$ and significantly improves memory efficiency, a crucial requirement when adapting to a large number of environments.
3.3 What’s in a context pool $\mathrm{P}$
The content of the context pool $\mathrm{P}$ not only defines the candidate trajectories we get from Equation 8, but also the speed and memory cost of the meta-training process. We outline 3 pool-filling strategies for selecting $p$ neighboring contexts:
  • Random-All (RA): all $p$ distinct contexts can be selected by randomly drawing their indices from $\operatorname{[\![}1,m\operatorname{]\!]}$. By repeatedly doing so, we maximize variance and reduce overfitting by ensuring the broadest form of self-modulation – since information can always (in the stochastic limit) flow from any environment into $e$.
  • Nearest-First (NF): only the $p$ closest contexts to $\xi^{e}$ are selected. Each index is drawn from the discrete uniform distribution $\mathcal{U}\{1,m\}$. This effective strategy encourages environments to form clusters while maintaining intra-cluster self-modulation. (Note, however, that if $p=1$, then $\mathrm{P}=\{e\}$ itself, and no self-modulation occurs.) In our experiments, we observe that this strategy is the most balanced with regards to the context pool size, InD and OoD performance.
  • Smallest-First (SF): the lowest contexts in $L^{1}$ norm are selected first. Since an environment with context close to 0 can be interpreted as an environment-agnostic feature like in Kirchmeyer et al. (2022); Blanke and Lelarge (2024), this strategy prioritizes the flow of information from that base or canonical environment to the one of interest $e$.
@@ 3.4 InD and OoD Testing
We distinguish two forms of testing. For In-Domain (InD) testing, the environments are the same as the ones used during training. In InD testing data, the underlying parameters of the dynamical system we aim to reconstruct are unchanged, and so the meta-learned context vectors are reused. Out-of-Distribution (OoD) testing considers environments encountered during adaptation. The data is from the same dynamical system, but defined by different parameter values (either interpolated or extrapolated).
In all forms of testing, only the main predicted trajectory corresponding to $j=e$ or $j=e^{\prime}$ is used in the MSE and MAPE metrics computation (see Section C.3 for definitions), thus returning to a standard Neural ODE. Other candidate trajectories are aggregated to provide a measure of uncertainty. The initial conditions that define the trajectories are always unseen, although their distribution never changes from meta-training to meta-testing.
@@ 4 Main Results
In this section, we evaluate the effectiveness of our proposed framework by investigating two main questions. (i) How good are NCFs at resolving (P1) and (P2) for interpolation tasks (Section 4.2)? (ii) How does our framework compare to state-of-the-art Meta-Learning baselines (Section 4.3)?
@@ 4.1 Experimental Setting
The NCF framework is evaluated on five seminal benchmarks. The Simple Pendulum (SP) models the periodic motion of a point mass suspended from a fixed point. The Lotka-Volterra (LV) system models the dynamics of an ecosystem in which two species interact. Additional ODEs include a simple model for yeast glycolysis: Glycolytic-Oscillator (GO) Daniels and Nemenman (2015), and the more advanced Sel’kov Model (SM) Strogatz (2018); Sel’Kov (1968). Like GO, SM is non-linearly parameterized, but additionally exhibits starkly different behaviors when its key parameter is varied (see LABEL:fig:attractors). Finally, we consider three PDEs, all with periodic boundary conditions, and cast as ODEs via the method of lines: the non-linear oscillatory Brusselator (BT) model for autocatalytic chemical reactions Prigogine and Lefever (1968), the Gray-Scott (GS) system also for reaction-diffusion in chemical settings Pearson (1993), and Navier-Stokes (NS) for incompressible fluid flow Stokes et al. (1851).
@@
For all problems, the parameters and initial states are sampled from distributions representative of real-world problems observed in the scientific community, and the trajectories are generated using a time-adaptive 4th-order Runge-Kutta solver Virtanen et al. (2020). For LV, GO, GS, and NS, we reproduce the original guidelines set in Kirchmeyer et al. (2022), while exposing the data for ODE and PDE problems alike via a common interface. Such use of synthetic data is a common practice in this emerging field of generalizable dynamical systems, where the search for unifying benchmarks remains an open problem Massaroli et al. (2020). This need for shared datasets and APIs has motivated the GEN-DYNAMICS:https://github.com/ddrous/gen-dynamics open-source initiative, our third and final contribution with this paper. Further details, along with the data generation process, are given in Appendix C.
@@
We now highlight a few key practical considerations shared across experiments. We use the 3-networks architecture depicted in Figure 2b to suitably process the state and context variables. The dimension of the context vector, the context pool’s size and filling strategy, and the numerical integration scheme vary across problems. For instance, we set $d_{\xi}=1024$ for SP and LV, $d_{\xi}=202$ for NS, and $d_{\xi}=256$ for all other problems; while $p=2$ for LV and SM, $p=4$ for LV and GO, and $p=3$ for all PDE problems. Other hyperparameters are carefully discussed in Appendix B.
@@
4.2 Interpolation Results
This experiment explores the SP problem discussed in Figure 1. During meta-training, we use 25 environments with the gravity $g$ regularly spaced in $[2,24]$. Each of these environments contains only 4 trajectories with the initial conditions $x^{e}_{i}(0)\sim\left(\mathcal{U}(-\frac{\pi}{3},\frac{\pi}{3}),\mathcal{U}(-1,1)\right)^{T}$. During adaptation, we interpolate to 2 new environments with $g\in\{10.25,14.75\}$.
Table 2: In-domain and adaptation testing MSEs ($\downarrow$) with OFA and NCF-$\mathit{t}_{1}$ on the SP problem.
In-Domain OoD
OFA (×$\times 10^{0}$) 1.14 $\pm$ 0.002 3.31 $\pm$ 0.026
NCF (×$\times 10^{-3}$) 4.82 $\pm$ 0.08 0.25 $\pm$ 0.02
Results in Table 2 emphasize the merits of Meta-Learning approaches like NCFs in lieu of naive baseline like One-For-All (OFA), where one context-agnostic vector field is learned for all environments indiscriminately. [^3] We refer the reader to Yin et al. (2021a) for a deeper discussion on One-For-All and One-Per-Env (one vector field for each environment). Additional results examining the learned context vectors are reported in Section C.2.
4.3 Extrapolation Results
A Meta-Learning algorithm is only as good as its ability to extrapolated to unseen environments. In this section we tackle several one-shot generalization problems, i.e. $S^{\prime}=1$. We consider the LV, GO, SM, GS, BT, and NS problems, which all involve adaptation environments outside their meta-training distributions. We compare both variants of NCF to two baselines: CAVIA Zintgraf et al. (2019) which is conceptually the closest GBML method to ours, and CoDA-$\ell_{1}$ Kirchmeyer et al. (2022). Their hyperparameters, laid out in Appendix B, are tuned for a balance of computational efficacy and performance, all the while respecting each baseline’s key assumptions.
At similar parameter counts [^4] This accounts for the total number of learnable parameters in the framework excluding the context vectors, whose size might vary with the method as explained in Appendix B., Table 1 shows that CAVIA is the least effective for learning all six physical systems, with a tendency to overfit on the few-shots it receives during both meta-training and meta-testing Mishra et al. (2017). NCF-$\mathit{t}_{2}$ achieves SoTA OoD results on 5 out of 6 problems. While CoDA retains its superiority on GS, we find that it struggles on non-linear problems, most notably on the SM problem whose trajectories go through 3 distinct attractors in a Hopf bifurcation. This shows that CoDA’s ability to automatically select low-rank adaptation subspaces via a linear hypernetwork decoder is limited, particularly for highly nonlinear systems. We note that NCF-$\mathit{t}_{1}$ equally fails to accurately resolve the SM problem, thus illustrating the value of second-order Taylor expansion and proximal alternating minimization lacking in its design but present in NCF-$\mathit{t}_{2}$. This suggests that proximal operators and Taylor-based regularization prevent NCF-$\mathit{t}_{2}$ from learning spurious associations, which is a problem commonly associated with poor OoD generalization Mouli et al. (2024).
@@ 4.3 Extrapolation Results
We report in Figure 3 how well NCF performs on 625 adaptation environments obtained by varying its parameters $\beta$ and $\delta$ on a 25$\times$25 uniform grid. It shows a consistently low MAPE below 2%, except in the bottom right corner where the MAPE rises to roughly 15%, still a remarkably low value for this problem. We highlight the 9 meta-training environments in yellow, and the 4 environments used for OoD adaptation for Table 1 in indigo. Remarkably, our training environment’s convex-hull where adaptation is particularly low is much larger compared to CoDA’s (Kirchmeyer et al., 2022, Figure 3).
@@
We investigate how the models perform across the SM attractors depicted in LABEL:fig:attractors: training environments $e_{1}$ and $e_{2}$ fall into a limit cycle (L1), $e_{3}$ and $e_{4}$ collapse to a stable equilibrium (E), while $e_{5}$ and $e_{6}$ fall into another limit cycle (L2). LABEL:fig:erropergroup presents the adaptation MSE for each environment in the OoD testing dataset. While CoDA equally fails to capture all 3 attractors, we observe that CAVIA and NCF-$\mathit{t}_{2}$ tend to favor E. But unlike CAVIA, NCF-$\mathit{t}_{2}$ succeeds in capturing the other limit cycles as well, as evidenced by the low MSE on all environments.
@@ 5.1 Benefits of NCFs
Neural Context Flows provide a robust but flexible framework for learning differential equations. They can handle irregularly-sampled sequences like time series, and can easily be extended to general regression tasks Finn et al. (2017). Some of their other desirable properties are highlighted below.
@@
Massively parallelizable.
Equation 10 indicates that NCFs are massively parallelizable along 3 directions. Indeed, evaluations of $\mathcal{L}$ can be vectorized across $m\times p$ environments, and across all $S$ trajectories. This leads to better use of computational resources for meta-training. We provide details on such vectorized NCF implementation in Appendix E, along with a thorough discussion on its scalability in Section A.6.
@@
Intrepretable.
Understanding how a model adapts to new physical settings is invaluable in many scientific scenarios. Moving information from one environment to another via context-affine transformations Blanke and Lelarge (2024) provides a powerful framework for explaining Multi-Task Learning. With contextual self-modulation via Taylor expansion, we generalize this framework while maintaining computational efficiency. In Figure 6 for instance, we leverage Proposition 1 of Blanke and Lelarge (2024) to showcase system identification with NCF-$\mathit{t}_{1}$, where the underlying physical parameters $\beta$ and $\delta$ of the LV problem are recovered up to a linear transform. The corresponding experiment is carefully detailled in Section A.2.
@@
The relative simplicity of our formulation makes NCFs easily adaptable to other Neural ODE models. For instance, it is straightforward to augment the vector field as in APHYNITY Yin et al. (2021b), augment the state space as in ANODE Dupont et al. (2019), or control the vector field as in Massaroli et al. (2020).
@@
Uncertainty requirements are increasingly critical in Meta-learned dynamical systems Liu et al. . NCFs provide, by their very definition, a measure of the uncertainty in the model. During testing, all candidate trajectories stemming from available contexts (meta-trained, adapted, or both) can be used to ascertain when or where the model is most uncertain in its predictions. In Figure 7 for instance, we visualize the means and standard deviations (scaled 10 folds for visual exposition) of 9 and 21 candidate forecasts for the Lotka-Volterra and Sel’kov Model problems, respectively.
@@ 5.2 Limitations
Theory.
While massively parallelizable, extendable, interpretable, and applicable to countless areas well outside the learning of dynamics models, Neural Context Flows still need additional theoretical analysis to explain their effectiveness. While the provided uncertainty estimation partly compensates for this shortage, we believe that this creates an interesting avenue, along with other limitations outlined below, for future work.
@@
Another limitation faced by Neural Context Flows lies within the assumptions they make. While differentiability of the vector field $f$ wrt $\xi$ is encountered with the majority of dynamical systems in science and engineering, some are bound to be fundamentally discontinuous. NCFs break down in such scenarios and would benefit from the vast body of research in numerical continuation Allgower and Georg (2012).
Hyperparameters.
Finally, NCF introduces several new hyperparameters such as the context pool size $p$, the context size $d_{\xi}$, the proximity parameter $\beta$ in Algorithm 1, and many more. Although we offer insights into their vital role in the Appendix D, we acknowledge that tuning them complicates the training process.
5.3 Broader Impact
While the benefits of NCFs are evidenced in Section 5.1, its negative impacts should not be neglected. For instance, malicious deployment of such adaptable models in scenarios they were not designed for could lead to serious adverse outcomes. This said, our data and models are openly available with minimal risk of misuse.

### v5

@@ 3.2 Adaptation (or Meta-Testing)
Our adaptation rule, as outlined in Algorithm 2 is extremely fast, converging in seconds for trainings that took hours. In scenarios where we want to adapt to more than one environment, we outline a bulk version in Section A.5 (see Algorithm 4). Although the bulk adaptation algorithm is parallelizable and framed in the same way as during meta-training, it does not allow for flow of contextual information, since this causes significant accuracy degradation. Most importantly, disabling the Taylor expansion at this stage limits the size of the context pool to $p=1$ and significantly improves memory efficiency, an important resource when adapting to a large number of environments.
@@ 3.3 InD and OoD Testing
We distinguish two forms of testing. For In-Domain (InD) testing, the environments are the same as the ones used during training. In InD testing data, the underlying parameters of the dynamical system we aim to reconstruct are unchanged, and so the meta-learned context vectors are reused. Out-of-Distribution (OoD) testing considers environments encountered during adaptation. The data is from the same dynamical system, but defined by different parameter values (either interpolated or extrapolated). In all forms of testing, only the main predicted trajectory corresponding to $j=e$ or $j=e^{\prime}$ is used in the MSE and MAPE metrics computation (see Section B.4 for definitions), thus returning to a standard Neural ODE. Other candidate trajectories are aggregated to ascertain the model’s uncertainty. The initial conditions that define the trajectories are always unseen, although their distribution never changes from meta-training to meta-testing.
@@ 4 Main Results
In this section, we evaluate the effectiveness of our proposed framework by investigating two main questions. (i) How good are NCFs at resolving (P1) and (P2) for interpolation tasks (Section 4.2)? (ii) How does our framework compare to SoTA Meta-Learning baselines (Section 4.3)? Further questions regarding interpretability, scalability, and uncertainty estimation are formulated and addressed in Sections A.2, B.2 and A.7 respectively.
@@ 4.1 Experimental Setting
The NCF framework is evaluated on five seminal benchmarks. The Simple Pendulum (SP) models the periodic motion of a point mass suspended from a fixed point. The Lotka-Volterra (LV) system models the dynamics of an ecosystem in which two species interact. Additional ODEs include a simple model for yeast glycolysis: Glycolytic-Oscillator (GO) Daniels & Nemenman (2015), and the more advanced Sel’kov Model (SM) Strogatz (2018); Sel’Kov (1968). Like GO, SM is non-linearly parameterized, but additionally exhibits starkly different behaviors when its key parameter is varied (see LABEL:fig:attractors). Finally, we consider three PDEs, all with periodic boundary conditions, and cast as ODEs via the method of lines: the non-linear oscillatory Brusselator (BT) model for autocatalytic chemical reactions Prigogine & Lefever (1968), the Gray-Scott (GS) system also for reaction-diffusion in chemical settings Pearson (1993), and Navier-Stokes (NS) for incompressible fluid flow Stokes et al. (1851).
@@
For all problems, the parameters and initial states are sampled from distributions representative of real-world problems observed in the scientific community, and the trajectories are generated using a time-adaptive 4th-order Runge-Kutta solver Virtanen et al. (2020). For LV, GO, GS, and NS, we reproduce the original guidelines set in Kirchmeyer et al. (2022), while exposing the data for ODE and PDE problems alike via a common interface. Such use of synthetic data is a common practice in this emerging field of generalizable dynamical systems, where the search for unifying benchmarks remains an open problem Massaroli et al. (2020). This need for shared datasets and APIs has motivated the Gen-Dynamics open-source initiative, our third and final contribution with this paper. Further details, along with the data generation process, are given in Appendix B.
@@
We now highlight a few key practical considerations shared across experiments. We use the 3-networks architecture depicted in Fig. 2b to suitably process the state and context variables. The dimension of the context vector, the context pool’s size and filling strategy, and the numerical integration scheme vary across problems. For instance, we set $d_{\xi}=1024$ for LV, $d_{\xi}=202$ for NS, and $d_{\xi}=256$ for all other problems; while $p=2$ for LV and SM, $p=4$ for LV and GO, and $p=3$ for all PDE problems. Other hyperparameters are carefully discussed in Appendix D.
4.2 Interpolation Results
Table 1: Training and adaptation testing MSEs ($\downarrow$) with OFA, OPE, and NCF-$\mathit{t}_{1}$ on the SP problem.
Train (×$\times 10^{-1}$) Adapt (×$\times 10^{-3}$)
OFA 9.49 $\pm$ 0.04 115000 $\pm$ 3200
OPE 0.18 $\pm$ 0.02 459.0 $\pm$ 345.0
NCF 0.10 $\pm$ 0.03 0.0356 $\pm$ 0.001
This experiment explores the SP problem discussed in Fig. 1. During meta-training, we use 25 environments with the gravity $g$ regularly spaced in $[2,24]$. Each of these environments contains only 4 trajectories with the initial conditions $x^{e}_{i}(0)\sim\left(\mathcal{U}(-\frac{\pi}{3},\frac{\pi}{3}),\mathcal{U}(-1,1)\right)^{T}$. During adaptation, we interpolate to 2 new environments with $g\in\{10.25,14.75\}$, each with 1 trajectory. For both training and adaptation testing scenarios, we generate 32 separate trajectories.
Table 1 emphasize the adaptation-time merits of Meta-Learning approaches like NCFs in lieu of baselines where one context-agnostic vector field is trained for all environments indiscriminately (One-For-All or OFA), or for each environment independently (One-Per-Env or OPE). Additional results and further analysis of these differences is done in Section B.3.
4.3 Extrapolation Results
A Meta-Learning algorithm is only as good as its ability to extrapolated to unseen environments. In this section we tackle several one-shot generalization problems, i.e. $S^{\prime}=1$. We consider the LV, GO, SM, GS, BT, and NS problems, which all involve adaptation environments outside their meta-training distributions. We compare both variants of NCF to two baselines: CAVIA Zintgraf et al. (2019) which is conceptually the closest GBML method to ours, and CoDA-$\ell_{1}$ Kirchmeyer et al. (2022). Their hyperparameters, laid out in Appendix D, are tuned for a balance of computational efficacy and performance, all the while respecting each baseline’s key assumptions.
At similar parameter counts [^3] This accounts for the total number of learnable parameters in the framework excluding the context vectors, whose size might vary with the method as explained in Appendix D., Table 2 shows that CAVIA is the least effective for learning all six physical systems, with a tendency to overfit on the few-shots it receives during both meta-training and meta-testing Mishra et al. (2017). NCF-$\mathit{t}_{2}$ achieves SoTA OoD results on 5 out of 6 problems. While CoDA retains its superiority on GS, we find that it struggles on non-linear problems, most notably on the SM problem whose trajectories go through 3 distinct attractors in a Hopf bifurcation.
This shows that CoDA’s ability to automatically select low-rank adaptation subspaces via a linear hypernetwork decoder is limited, particularly for highly nonlinear systems. We note that NCF-$\mathit{t}_{1}$ equally fails to accurately resolve SM and other non-linear problems, thus illustrating the value of second-order Taylor expansion. This suggests that higher-order Taylor-based regularization prevents NCF-$\mathit{t}_{2}$ from learning spurious associations, which is a problem commonly associated with poor OoD generalization Mouli et al. (2024).
@@ 4.3 Extrapolation Results
We report in Fig. 3 how well NCF performs on 625 adaptation environments obtained by varying its parameters $\beta$ and $\delta$ on a 25$\times$25 uniform grid. It shows a consistently low MAPE below 2%, except in the bottom right corner where the MAPE rises to roughly 15%, still a remarkably low value for this problem. We highlight the 9 meta-training environments in yellow, and the 4 environments used for OoD adaptation for Table 2 in indigo. Remarkably, our training environment’s convex-hull where adaptation is particularly low is much larger compared to CoDA’s (Kirchmeyer et al., 2022, Figure 3).
@@
We investigate how the models perform across the SM attractors depicted in LABEL:fig:attractors: training environments $e_{1}$ and $e_{2}$ fall into a limit cycle (L1), $e_{3}$ and $e_{4}$ collapse to a stable equilibrium (E), while $e_{5}$ and $e_{6}$ fall into another limit cycle (L2). LABEL:fig:erropergroup presents the adaptation MSE for each environment in the OoD testing dataset. While CoDA equally fails to capture all 3 attractors, we observe that CAVIA and NCF-$\mathit{t}_{2}$ tend to favor E. But unlike CAVIA, NCF-$\mathit{t}_{2}$ succeeds in capturing limit cycles as well, as evidenced by the low MSE on all environments.
@@ 5.1 Benefits of NCFs
Neural Context Flows provide a powerful and flexible framework for learning differential equations. They can handle irregularly-sampled sequences like time series, and can easily be extended to general regression tasks Finn et al. (2017). Some of their other desirable properties are highlighted below.
Massively parallelizable.
Eq. 10 indicates that NCFs are massively parallelizable along 3 directions. Indeed, evaluations of $\mathcal{L}$ can be vectorized across $m\times p$ environments, and across all $S$ trajectories. This leads to better use of computational resources for meta-training. We provide details on such vectorized NCF implementation in Appendix E, along with a thorough discussion on its scalability in Section A.7.
@@
Interpretable.
Understanding how a model adapts to new physical settings is invaluable in many scientific scenarios. Moving information from one environment to another via context-affine transformations provides a powerful framework for explaining Multi-Task Learning Blanke & Lelarge (2024). With contextual self-modulation via Taylor expansion, we generalize this framework while maintaining computational efficiency. In Fig. 6 for instance, we showcase system identification with NCF-$\mathit{t}_{1}$, where the underlying physical parameters $\beta$ and $\delta$ of the LV problem are recovered up to a linear transform. The corresponding experiment is detailed in Section A.2 which further illustrates NCF’s robustness to noise during system identification.
Proposition 2 (Identifiability of affine systems).
Assume $d_{\xi}\geq d_{c}$, that $P$ is full-rank, and that $f_{\textup{true}}$ is differentiable in its second argument. In the limit of zero training loss in Eq. 10, $f_{\theta}$ trained with first-order Taylor expansion and the Random-All pool-filling strategy [^4] A discussion on which neighboring contexts should go into the context pool $\mathrm{P}$ is provided in Section A.6. is affine on an open region of $\mathbb{R}^{d_{\xi}}$. Furthermore, there exists $Q\in\mathbb{R}^{d_{c}\times d_{\xi}}$ and $q\in\mathbb{R}^{d_{c}}$ such that for any meta-trained $\xi\in\{\xi^{e}\}_{e=1}^{m}$ and its corresponding underlying parameter $c\in\{c^{e}\}_{e=1}^{m}$, we have $c=Q\xi+q$.
To theoretically demonstrate the capacity to identify physical systems, we consider a learnable vector field $f_{\text{true}}:\mathbb{R}^{d}\times\mathbb{R}^{d_{c}}\rightarrow\mathbb{R}^{d}$ is that is affine i.e. $\exists P\in\mathbb{R}^{d\times d_{c}}\text{ and }p\in\mathbb{R}^{d}$ such that $f_{\text{true}}(\cdot,c)=Pc+p$. Under these assumptions, the predictor $f_{\theta}:\mathbb{R}^{d}\times\mathbb{R}^{d_{\xi}}\rightarrow\mathbb{R}^{d}$ satisfies Proposition 2 above, inspired by Proposition 1 of Blanke & Lelarge (2024). A detailed proof is provided in Section A.2.
@@
The relative simplicity of our formulation makes NCFs easily adaptable to other works involving Neural ODE models. For instance, it is straightforward to augment the vector field as in APHYNITY Yin et al. (2021b), augment the state space as in ANODE Dupont et al. (2019), or control the vector field as in Massaroli et al. (2020).
@@
Uncertainty requirements are increasingly critical in Meta-Learned dynamical systems Liu et al. . NCFs provide, by their very definition, a measure of the uncertainty in the model. During testing, all candidate trajectories stemming from available contexts (meta-trained, adapted, or both) can be used to ascertain when or where the model is most uncertain in its predictions. In Fig. 7 for instance, we visualize the means and standard deviations (scaled 10 folds for visual exposition) across 9 and 21 candidate forecasts for the Lotka-Volterra and Sel’kov Model problems, respectively. Additional results with quantitative metrics are provided in Section B.2.
@@ 5.2 Limitations
Further theory.
While massively parallelizable, extendable, interpretable, and applicable to countless areas well outside the learning of dynamics models, Neural Context Flows still need additional theoretical analysis to explain their effectiveness and uncover failure modes. While the provided Propositions 1 and 2 partly compensates for this shortage, we believe that this creates an interesting avenue, along with other limitations outlined below, for future work.
@@
Another limitation faced by Neural Context Flows lies within the assumptions they make. While differentiability of the vector field $f$ wrt $\xi$ is encountered with the majority of dynamical systems in science and engineering, some are bound to be fundamentally discontinuous. NCFs break down in such scenarios and would benefit from the vast body of research in numerical continuation Allgower & Georg (2012).
Additional hyperparameters.
Finally, NCF introduces several new hyperparameters such as the context size $d_{\xi}$, the context pool size $p$, the pool-filling strategy, the proximity coefficient $\beta$ in Algorithm 1, and many more. Although we offer insights into their roles in the Appendix C, we acknowledge that tuning them complicates the training process.